# An Entropy-based Text Watermarking Detection Method

### v1

@@ Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token should be adjusted according to its entropy during watermark detection, rather than setting the weight of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data will be available online.
@@ 1 Introduction
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the tokens, it leads the generated text to contain mostly watermarked (green) tokens. In the watermark detection process, the degree of watermark is usually judged by detecting the number of watermarked tokens in the text.
@@
Although the current text watermarking algorithm performs well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when modified.
@@
In figure 1, we illustrate the watermark ratio in tokens of different entropy levels. A token with higher entropy will have a significantly larger probability to be modified and watermarked. On the contrary, a low-entropy token has a much lower probability.
@@
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the state of these unmodified low-entropy tokens. For example, the low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed an entropy-based selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold, thus excluding tokens that are difficult to be modified by the watermark generator. However, this method also has some problems. First, their method requires a manually set entropy threshold, which needs to be determined by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
@@
To address the aforementioned problems on low-entropy watermark detection, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the importance of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Therefore, we assign an importance weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm failed to watermark a low-entropy token, it will make minimal impact to the detected watermark level. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy level. There is no need to switch between different detectors facing various kinds of input texts.
@@
In the experiment, we evaluated the detection accuracy of our proposed methods in various low-entropy generation tasks, including code generation. We also test the versatility of our method in text of high entropy level, such as document-to-text generation. Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy.
@@
  • We provide a theoretical analysis of the detection accuracy in our method in a low-entropy setting.
  • Through experiments, we show that our method not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
@@ 2 Related Work
Watermark message can be embedded into existing machine-generated text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al. (2006); Yang et al. (2022); Yoo et al. (2023a); Yang et al. (2023). Those methods typically replace words with their alternatives to place watermark. For example, Topkara et al. (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words. To ensure that the repalced words would cause minimal changes to the text semantics, context-aware lexical substitution was introduced. For example, Yang et al. (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning. To improve the robustness of the algorithm against possible corruptions, Yoo et al. (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin. However, the level of modification to an existing text is often limited due to the need to protect text quality, so the capacity and effectiveness of lexical-based watermarks are largely restricted.
@@
Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits. Christ et al. (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens. Detection process is to check the correlation of the input text with random number sequences. To enhance resistance to text modifications, Kuditipudi et al. (2023) utilized the Levenshtein distance in matching the text with number sequences. For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al. (2023a) based on previous tokens during text generation. Specifically, the model vocabulary will be separated into green and red lists during each generation step. And the logit value of every green token will be enhanced by a positive constant. There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al. (2023b); Wang et al. (2023), robustness Zhao et al. (2023); Liu et al. (2023b); Kirchenbauer et al. (2023b); Ren et al. (2023) and unforgeabilityLiu et al. (2023a); Fairoze et al. (2023). For example, Yoo et al. (2023b) extends the method to a multi-bit setting to allow the embedding of longer messages. Wang et al. (2023) proposed to utilize a proxy LLM for multi-bit message embedding. To increase robustness against editing attacks, Zhao et al. (2023) proposed to set a universal green and red lists during the text generation process. Liu et al. (2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green and red lists. Also, to enhance the unforgeability of text watermark, Liu et al. (2023a) introduced a neural network for text generation and detection. In this work, we focus more on the detection of low-entropy texts. Lee et al. (2023) designed a selective watermark generation and detection approach for low-entropy code generation. However, it requires a human code dataset to help manually determine the entropy threshold.
@@ 3.1 LLM
Here we introduce the necessary concepts used in this paper. A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response. Specifically, we denote the initial prompt as $x^{prompt}$. Then, at the $l$-th step, the input to the LLM is the $x^{prompt}$ and the already-generated tokens $T_{:(l-1)}$. The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input. The token sampling is based on the logits and specified sampling strategies, such as greedy sampling.
@@ 4 Method
In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and previously-proposed detection methods, then provide a detailed introduction to the proposed EWD algorithm.
@@ 4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this watermark embedding process would very likely produce a green-list (watermarked) token. An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability. As a result, the number of watermarked tokens and the consequence z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
@@ 4.2 Proposed Method
As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy. And we take a step further than SWEET: assigning an entropy-based weight to each token, representing their significance in the detection process.
In figure 1, we can observe that a high-entropy token has a larger probability of being watermarked or is easier to be modified and watermarked than a low-entropy token. Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited impact on the overall watermark level. Specifically, we argue that when an easy-to-watermark token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token. That is to say, The impact of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the impact of a watermarked token to yielding a positive detection result:
$I(Detect(T)=\textbf{1}(t\in G)|t)\propto Entropy(t).$ (3)
When reflected in the real world, this impact is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all watermarked token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a non-watermarked low-entropy token than a non-watermarked high-entropy token, and a larger gain when observing a watermarked high-entropy token than a watermarked low-entropy token.
@@
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the importance weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights:
$W=Linear(Norm(SE)).$ (5)
After obtaining the weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the watermarked tokens and calculate the z-statistic using Eq.4 given the sum and weight vector of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, such as code generation, there are a large number of hard-to-modify tokens. When detecting watermarked texts with our proposed EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing non-watermarked texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
@@ 5 Theoretical Analysis
In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous methods, KGWKirchenbauer et al. (2023a) and SWEETLee et al. (2023).
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. Therefore, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
Table 1: Main results on different datasets using KGW, SWEET and our method for detection.
@@ 5.2 Type-II Error
Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written.
For the purposes of analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and that the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated new scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being watermarked by a watermark generator:
@@
KGW. Based on Eq.6, the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%, or in other words, the detection rate of watermarked text is less than 15.4%. This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
SWEET. Following the detection scheme in SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection. The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively. Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%, or in other words, the detection rate of watermarked text is around 58%.
EWD. We assigned higher weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy. Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant. Based on Eq.6, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3%, or in other words, the detection rate of watermarked text is around 66.7% (Proof and details in Appendix A). Overall, Our method has a type-II error of 33.3%, much lower than KGW and SWEET. Empirical experiments also demonstrate similar results.

### v2

@@ Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be adjusted according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data will be available online.
@@ 1 Introduction
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the token distributions, it leads the generated text to contain mostly watermarked (green) tokens. In the watermark detection process, the degree of watermark is usually judged by the number of watermarked tokens in the text.
@@
Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when forcefully modified Tu et al. (2023).
@@
In figure 1, we illustrate the watermark ratio in tokens of three different entropy levels. A token with higher entropy will have a significantly larger probability to be modified and watermarked. On the contrary, it is rare to see a green token in low-entropy tokens.
@@
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the states of these unmodified low-entropy tokens even when the watermark generator managed to modify all high-entropy tokens. For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed a selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold. However, this method has some problems. First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
@@
To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the influence of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Therefore, we assign an influence weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm failed to watermark a low-entropy token, it will make minimal impact to the detected watermark level. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels. There is no need to switch between different detectors facing various kinds of input texts.
@@
In the experiment, we evaluated the detection accuracy of EWD in low-entropy code generation task. We also test the versatility of our method in text of high entropy levels, such as document-to-text generation. Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy. Besides, we included evaluation of our proposed EWD against watermark-removal attacks.
@@
  • We provide a theoretical analysis of the detection accuracy in a low-entropy setting.
  • We showed that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
@@ 2 Related Work
Watermark message can be embedded into existing text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al. (2006); Yang et al. (2022); Yoo et al. (2023a); Yang et al. (2023). Those methods typically replace words with their alternatives to place watermark. For example, Topkara et al. (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words. To ensure that the repalced words would cause minimal changes to the text semantics, Yang et al. (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning. To improve the robustness of the algorithm against possible corruptions, Yoo et al. (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin. However, the level of modification to an existing text is often limited to protect text quality, thus restricting the capacity and effectiveness of lexical-based watermarks.
@@
Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits. Christ et al. (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens. To enhance resistance to text modifications, Kuditipudi et al. (2023) utilized the Levenshtein distance in matching the text and numbers. For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al. (2023a) based on previous tokens during text generation. There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al. (2023b); Wang et al. (2023), robustness Zhao et al. (2023); Liu et al. (2023b); Kirchenbauer et al. (2023b); Ren et al. (2023) and unforgeabilityLiu et al. (2023a); Fairoze et al. (2023). For example, Yoo et al. (2023b) extends the method to a multi-bit setting.Wang et al. (2023) proposed to utilize a proxy LLM for multi-bit message embedding. To increase robustness against editing attacks, Zhao et al. (2023) proposed to set a universal green list during the generation process. Liu et al. (2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green lists. Also, to enhance the unforgeability of text watermark, Liu et al. (2023a) introduced a neural network for text generation and detection. In this work, we focus only on the detection of low-entropy texts. Lee et al. (2023) designed a selective watermark generation and detection approach for low-entropy code generation. However, it requires a human code dataset to help manually determine the entropy threshold.
@@ 3.1 LLM
Here we introduce the necessary concepts in this paper. A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response. Specifically, we denote the initial prompt as $x^{prompt}$. Then, at the $l$-th step, the input to the LLM is the $x^{prompt}$ and the already-generated tokens $T_{:(l-1)}$. The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input.
@@ 4 Method
In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and traditional detection methods, then provide a detailed introduction to the proposed EWD.
@@ 4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list (watermarked) token. An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability. As a result, the number of watermarked tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1. This is because to the detector, tokens are of the same influence, no matter how difficult it is to modify them.
@@ 4.2 Proposed Method
As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy. And we took a step further than SWEET: assigning an entropy-based weight to each token, representing their influence in the detection process.
In figure 1, we can observe that a high-entropy token is easier to be modified and thus has a larger probability of being watermarked than a low-entropy token. Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited influence on the overall watermark level. Specifically, we argued that when an easy-to-modify token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token. That is to say, The influence of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the influence of a watermarked token to yielding a positive detection result:
$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t).$ (3)
When reflected in the real world, this influence index is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger influence weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all watermarked token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a non-watermarked low-entropy token than a non-watermarked high-entropy token, and a larger gain when observing a watermarked high-entropy token than a watermarked low-entropy token.
@@
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights. After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the watermarked tokens and calculate the z-statistic using Eq.4 given the sum and weight vector of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, With our EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing human-written texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
@@ 5 Theoretical Analysis
In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous detection methods, KGW and SWEET.
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. While Zhao et al. (2023) set window size in the watermark generator as zero and exhibited major token dependence, $i.e.$, the same tokens in the text are of same watermark status (either all green or all red), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened. For example, the same token with different prefix would exhibit different watermark status. Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the watermark status of each token is independent. Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
Table 1: Main results on different datasets using KGW, SWEET and our EWD for detection.
@@ 5.2 Type-II Error
Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written. For analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being watermarked by a watermark generator:
@@
KGW. Based on Eq.5, the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%. This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
SWEET. Following SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection. The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively. Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%.
EWD. We assigned higher influence weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy. Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant. Based on Eq.5, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3% (Proof and details in Appendix A). Overall, Our method has a type-II error much lower than KGW and SWEET.